OpenStudy (anonymous):

The two legs of a right triangle are 8 inches and 3 inches. Find the hypotenuse of the triangle.

OpenStudy (anonymous):

= sqrt(8^2 + 3^2) = sqrt73 =8.6

OpenStudy (anonymous):

8.6

OpenStudy (anonymous):

\[a^2=b^2+c^2\] so hyp = \[\sqrt{64+9}\] = 8.5440

OpenStudy (anonymous):

so rounded the answer is 8.5 surely

OpenStudy (anonymous):

yes u r right key

OpenStudy (anonymous):

when u r happy with the help acura, plz click good answer at whoever u want, to close this post